OpenStudy (anonymous):

Given the parent functions f(x) = 5x − 1 and g(x) = 3x − 9, what is g(x) − f(x)?

OpenStudy (usukidoll):

given: g(x) = 3x-9 f(x) = 5x-1 and we need to solve g(x) - f(x)

OpenStudy (usukidoll):

so we plug these functions into g(x)-f(x) 3x-9-(5x-1) now distribute the negative on the right hand side of the equation

OpenStudy (anonymous):

3x-9(-5x+1)?

OpenStudy (usukidoll):

ok... that's good 3x-9-5x+1 now we combine like terms.. we can also rearrange this equation to make it easier to solve so instead of 3x-9-5x+1 we can rewrite it as 3x-5x-9+1

OpenStudy (anonymous):

oh the equation is 3^x

OpenStudy (anonymous):

im sorry

OpenStudy (usukidoll):

ahhhhhhhh! ok no problem... just rewrite what f(x) and g(x) was supposed to be

OpenStudy (usukidoll):

ok was g(x) = 3^x-9 ?

OpenStudy (anonymous):

f(x)=5x-1 g(x)=3^x-9

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (usukidoll):

ok we do the same process again g(x) -f(x) 3^x-9-(5x-1)

OpenStudy (usukidoll):

actually we have already done the distribution of the negative, 3^x-9-5x+1 now combine like terms for -9+1

OpenStudy (anonymous):

-8

OpenStudy (usukidoll):

\[3^x-9+1-5x\]

OpenStudy (usukidoll):

yup \[3^x-8-5x\]

OpenStudy (usukidoll):

since we can't combine any more terms, that's the answer

OpenStudy (anonymous):

Given the parent functions f(x) = log10 x and g(x) = 5x − 2, what is f(x) • g(x)? f(x) • g(x) = log10 x5x − 2 f(x) • g(x) = log10 (5x − 2)x f(x) • g(x) = 5x log10 x + 2 log10 x f(x) • g(x) = 2 log10 x − 5x log10 x

OpenStudy (anonymous):

let me type the equation correctly

OpenStudy (anonymous):

\[f(x)=\log_{10} x \] g(x)=5x-2

OpenStudy (anonymous):

firt answer is log10 x^5-2 second is log10(5x-2)^x

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (usukidoll):

we are using multiplication \[f(x) \cdot g(x) \] once again we place our f(x) function and g(x) into the formula

OpenStudy (usukidoll):

\[\log10_x[5x-2]\] use distribution

OpenStudy (anonymous):

what do you distribute

OpenStudy (usukidoll):

distribute the log

OpenStudy (usukidoll):

\[\log_{10}x[5x-2]\]

OpenStudy (anonymous):

i dont have a graphing or scientific calc to solve

OpenStudy (usukidoll):

we can just distribute the log